OpenStudy (anonymous):

Meg rowed her boat upstream a distance of 64 miles and then rowed back to the starting point. The total time of her trip was 20 hours. If the rate of the current was 6 mph, find the average speed of the boat relative to the water.

OpenStudy (anonymous):

let the speed of the boat in still water = x mph then speed while going downstream = (x+6) mph and speed while going upstream = (x - 6) mph distance = speed x time => time = distance/speed time1 while going downstream = 64/(x + 6) hr time2 while going upstream = 64/(x - 6) hr given that time1 + time2 = 20 => 64/(x + 6) + 64(x - 6) = 20

OpenStudy (anonymous):

We have: 64/(b-6)+64/(b+6)=20 where b is the speed of the boat relative to the water. So (64(b+6)+64(b-6))/(b+6)(b-6)=20 128b=20(\[b ^{2}-36\]) 20b^2-128b-720=0 or 5b^2-32b-180=0 (5b+18)(b-10)=0 We throw out the negative one, giving b=10 mph You can plug it back in to check.

OpenStudy (anonymous):

thank you so much 10 is the answer. :-)

OpenStudy (anonymous):

goblue has given the further steps.. however if u want me to complete my solution, let me know...

OpenStudy (radar):

If you do decide to complete Harikirat, you will need to correct the last line, it appears you exchange your division "/" to a multiplication. "=> 64/(x + 6) + 64(x - 6) = 20"

OpenStudy (anonymous):

thanks for pointing out the error :))

OpenStudy (radar):

I tried it, but got tangled up with the quadratic........I didn't discern the factors lol